9. Illustrate a downward shift in the IS curve. Describe three factors that could shift the IS-curve.

10.
a. What is the money multiplier?
b. How do we derive the money multiplier?
c. What is the magnitude in the money multiplier?
d. What accounts for this magnitude?
1
Expert's answer
2021-10-18T17:47:58-0400

(9)



A negative shock to investment shifts IS curve downwards reducing output, employment and interest rate in the short run.

The factors that could shift the IS curve include:

  • Increase in taxes- this shifts the IS curve downwards because it results in a decrease in the amount of savings and hence a fall in the level of investment.
  • Increase in government expenditure- this reduces the level of income and thus shifts the aggregate expenditure curve downwards.
  • Reduction in demand for money- the consequence of this is a fall in the toe rate of interest.

(10)

(a)

Money multiplier refers to a phenomenon of creating money in the economy in the form of creation of credit.

(b)

If we link money supply (M) to the monetary base(MB) and let m be the money multiplier;

"M=m\times MB" .

Deriving money multiplier I,

Assume the derived level of currency ,C and excess reserves, ER grow proportionally with checkable deposits D.

Then,

"c=\frac{C}{D}" = Currency ratio.

"e=\frac{ER}{D}" = Excess reserves ratio.

(c)

Magnitude of money multiplier is given by change in total money supply divided by change in monetary base (reserves).

(d)

The Reserve ratio which is the percentage of deposits that banks keep in liquid reserves account for the money multiplier magnitude.